# Is Repeated Bayesian Interim Analysis Consequence-Free?

### v1

@@ 3.1 Simulation setting
We conducted extensive simulation studies to evaluate the impact of repeated Bayesian interim analyses on various design metrics. The maximum sample size was set to $N=100$, and an equivalence margin of $\delta=0$ was used to reflect the setting of typical superiority trials. The data-generating prior, $\pi_{g}(\theta)$, was assumed to belong to the same probability distribution family as the user-specified prior, $\pi(\theta)$, but with potentially different hyperparameter values. For each outcome type and each scenario defined by $\pi(\theta)$, we simulated 50,000 trials. In the $k$th simulated trial, we first sampled $\theta^{k}$ from $\pi_{g}(\theta)$, then generated interim data $X^{k}_{1},X^{k}_{2},\cdots,X^{k}_{n}$ from $f(X_{i}|\theta^{k})$, e.g., $\mbox{Bernoulli}(\theta^{k})$ for binary endpoints or $N(\theta^{k},\sigma^{2})$ for normal endpoints. We implemented two interim analyses at the 40th and 70th patients, and a final analysis at the 100th patient, i.e., for n=40, 70n=40,\ 70, and $100$, the trial was stopped and $H_{0}$ rejected if $\mbox{Pr}(H_{a}|{\cal D}^{k}_{n})>C$, where ${\cal D}^{k}_{n}=\{X^{k}_{1},\cdots,X^{k}_{n}\}$.
@@
As a benchmark, we also evaluated the same design metrics for trials conducted without interim analyses. In these trials, the decision to reject $H_{0}$ was made at the end of the trial based on the criterion $\mbox{Pr}(H_{a}|{\cal D}^{k}_{N})>C$. The probability cutoff $C$ was calibrated to control the pFDR at 5% under the scenario where $\pi(\theta)=\pi_{g}(\theta)$. This same cutoff was then applied to designs with and without interim analyses; therefore, any differences observed can be directly attributed to the inclusion of interim analyses. It is important to note that we did not re-calibrate the decision cutoff $C$ for each individual scenario. As a result, in scenarios where $\pi(\theta)\neq\pi_{g}(\theta)$, the pFDR and other metrics for the fixed design may deviate from the nominal level of 0.05. However, this does not pose a concern for our investigation, as the primary objective is to compare results with and without interim analyses to assess their impact on operating characteristics, rather than to evaluate the absolute operating characteristics of a design across different scenarios.
@@
Below are the simulation parameters for single-arm trials; those for RCTs are provided in Supplementary Materials. For binary endpoint, we set the response rate of the (historical) control as $\theta_{0}=0.6$ and assumed the true response rate of the experimental arm generated from $\pi_{g}(\theta)=$ Beta(3, 3). We evaluated three classes of user-specified prior $\pi(\theta)$: (i) neutral priors centered at $\theta_{0}=0.6$; (ii) pessimistic priors centered at $0.5$; and (iii) optimistic priors centered at $0.7$. Within each class, we examined three priors with increasing spread (i.e., decreasing informativeness). Specifically, for $\pi(\theta)$ centered at 0.6, we considered Beta(18, 12), Beta(3.6, 2.4), and Beta(0.06, 0.04). For $\pi(\theta)$ centered at 0.5, we considered Beta(15, 15), Beta(3, 3), and Beta(0.05, 0.05). Notably, Beta(3, 3) matches $\pi_{g}(\theta)$. And for $\pi(\theta)$ centered at 0.7, we considered Beta(21, 9), Beta(4.2, 1.8), and Beta(0.07, 0.03). The probability cutoff was calibrated as previously described and set to $C=0.689$. Due to the discrete nature of the observed data, it was not always possible to find a cutoff yielding an exact 5% pFDR. This value of $C$ was selected to achieve an pFDR closest to, but not exceeding, 5%.
@@
For the normal outcome with $\sigma^{2}=1$, we assumed $\pi_{g}(\theta)=N(0,1)$ and set $\theta_{0}=0.25$. The neutral priors $\pi(\theta)$ centered at 0.25 included $N(0.25,0.2),N(0.25,1)$, and $N(0.25,1000)$. Optimistic and pessimistic priors were centered at 0 and 0.5, respectively. The probability cutoff was calibrated and set to $C=0.4$. For the normal outcome with unknown variance, we used the conjugate prior: $\pi_{g}(\theta|\sigma^{2})=N(0,\sigma^{2}/5)$ and $\pi_{g}(\sigma^{2})=\mbox{Inv-}\chi^{2}(5,40)$, with $\theta_{0}=0.25$. Priors $\pi(\theta,\sigma^{2})$ were constructed with $\theta$ centered at 0.25 (neutral), 0 (pessimistic), and 0.5 (optimistic), as detailed in Table 4. The probability cutoff was set at $C=0.63$.
@@
For the time-to-event outcomes, we assumed a 12-month follow-up period with patient accrual following a Poisson process at a rate of six patients per month. We set $\theta_{0}=8$ and used $\pi_{g}(\theta)=\mbox{Gamma}(12,2)$ and $\pi_{g}(\kappa)=\mbox{Gamma}(8,2)$. We evaluated three $\pi(\theta,\kappa)$ priors with $\theta$ centered at 8 (neutral), 6 (pessimistic), and 10 (optimistic), each with varying precision (see Table 5). The cutoff was calibrated and set to $C=0.61$ to control the pFDR at 5% in the absence of an interim analysis.
@@
The performance metrics were calculated as follows. pFDR was calculated as the proportion of trials in which $H_{0}$ was true among those in which $H_{0}$ was rejected, across the 50,000 simulated trials. FDR was then obtained by multiplying the pFDR by the proportion of trials in which $H_{0}$ was rejected. Bias was determined as the average difference $\hat{\theta}^{k}-\theta^{k}$, and MSE was computed as the mean of $(\hat{\theta}^{k}-\theta^{k})^{2}$, where $\hat{\theta}^{k}$ denotes the posterior mean of $\theta$ in the $k$th simulated trial. Coverage was defined as the percentage of trials in which $\theta^{k}$ fell within the 95% posterior credible interval for $\theta$ in the $k$th trial. Type I error rate A was computed as the percentage of rejected $H_{0}$ among those in which $H_{0}$ was true, and power was computed as the percentage of rejected $H_{0}$ among those in which $H_{0}$ was false. Type I error rate B was defined as one minus the coverage probability for the one-sided posterior credible interval when $H_{0}$ was true. Additionally, we report the average sample size across all simulated trials.
@@ 3.2 Results
Table 2 presents results for single-arm trials with a binary endpoint, using the design without interim analyses as the reference. Unless otherwise specified, the following discussion focuses on comparing designs with interim analyses to those without, in order to assess the impact of conducting multiple interim looks. Our findings are consistent with the theoretical investigation in Section 2.3, indicating that interim analyses can have a substantial effect on the operating characteristics of the design.
In the ideal scenario where the user-specified prior matches the data-generating prior (i.e., $\pi(\theta)=\pi_{g}(\theta)=$Beta(3, 3)), adding interim analyses does not affect the coverage rates of either one-sided or symmetric credible intervals, or the bias. However, incorporating interim analyses does lead to inflation in pFDR, type I error A and B, power, and MSE. For example, incorporating interim analyses increased the pFDR from 0.046 to 0.095 and type I error A from 0.018 to 0.041. This differential impact across various metrics may help explain discrepancies in the literature—while some authors have claimed that interim analyses and early stopping have no effect on the operating characteristics of a design, others have argued that they do. Although both Type I error rates A and B are inflated, the inflation of the latter is substantially smaller than that of the former, suggesting that Type I error B (the Type I error rate based on credible intervals) is more robust. This robustness stems from the fact that Type I error rate B partially inherits the properties of credible intervals, which are well calibrated and remain unaffected by interim analyses when $\pi(\theta)=\pi_{g}(\theta)$.
@@
Tables 3 and 4 present analogous results for normal endpoints with known and unknown variances, and Table 5 shows findings for time-to-event endpoints. Across all endpoint types, patterns were broadly consistent with those described above for the binary case. FDR values for these outcome types are reported in Tables S1 and S2 of the Supplementary Materials, showing similar patterns to those observed for pFDR. Results for RCTs are generally similar to those of single-arm trials; see Tables S3-S6 in Supplementary Materials for more details.
@@ 3.3 Sensitivity analyses
We examined the effect of altering the frequency of interim analyses. For binary and normal outcomes, where patient-level data are typically observed promptly, we employed a fully sequential design in place of the original two interim analyses at the 40th and 70th patients. In this design, early stopping was assessed after each patient beginning with the 20th, i.e., for $n=20,\cdots,N$, the trial was stopped and $H_{0}$ rejected if $\mbox{Pr}(H_{a}|{\cal D}^{k}_{n})>C$, where ${\cal D}^{k}_{n}=\{X^{k}_{1},\cdots,X^{k}_{n}\}$. For time-to-event endpoints, where outcomes are event-driven rather than patient-driven, interim analyses were conducted every 5 patients starting from the 20th, to allow time for meaningful accumulation of event information. Results for each outcome type under these settings are shown in Tables S7-S9 and were broadly consistent with those observed under the original design.
We examined the effect of altering the equivalence margin $\delta$. Specifically, we evaluated scenarios with positive values of $\delta$, meaning that the experimental arm must demonstrate a minimum improvement of $\delta$ (e.g., a minimal clinically important difference) over the control to be considered successful. Results for each outcome type under these settings are presented in Tables S10 - S12. The observed performance metrics followed patterns consistent with those under the original setting.
We further evaluated the performance metrics under alternative values of the probability of $H_{0}$ under $\pi_{g}(\theta)$, denoted by $\mbox{Pr}_{g}(H_{0})$. For example, for the normal outcome with known variance equal to 1 and equivalence margin $\delta=0$, the prior $\pi_{g}(\theta)=N(0,1)$ yields $\mbox{Pr}_{g}(H_{0})=0.6$. We also considered a more concentrated prior, $\pi_{g}(\theta)=N(0,0.09)$, which increases $\mbox{Pr}_{g}(H_{0})$ to 0.8. The corresponding results for continuous, binary, and time-to-event outcomes in single-arm trials, presented in Tables S13 - S15 of the Supplementary Materials, exhibited patterns consistent with those observed under the original prior specification.
For the normal outcome with unknown variance, the scale parameter $\sigma_{0}^{2}$ in $\pi(\sigma^{2})$ was originally set to 40, matching that in $\pi_{g}(\sigma^{2})$. In sensitivity analyses, we varied $\sigma^{2}_{0}$ to 20 and 60. The resulting metrics, shown in Tables S16 and S17, remained consistent with those from the main setting, further supporting the robustness of the findings.
@@ 4 Conclusion
We have presented a comprehensive investigation into the effects of repeated Bayesian interim analyses on various design metrics, including bias, MSE, the coverage probability of credible intervals, FDR, Type I error rates, and power. Our theoretical results establish, under general conditions and without reliance on any specific probability model, that repeated Bayesian interim analyses are not consequence free: they can systematically impact these inferential quantities. For some metrics (e.g., FDR), distortions arise even when the user-specified prior exactly matches the data-generating prior, and can be further exacerbated under prior misspecification. These theoretical insights are validated by extensive simulation studies, which confirm that repeated Bayesian interim analyses can affect operating characteristics of trials, particularly when vague or noninformative priors are used. This is of great concern as, in practice, noninformative or vague priors are most likely to be used.
@@
Data Availability
Data sharing is not applicable in this paper as all data in this paper are computer simulated.

### v2

@@ 3.1 Simulation setting
We conducted simulation studies to evaluate the impact of repeated Bayesian interim analyses on various design performance metrics. The maximum sample size was set to $N=100$. The data-generating prior, $\pi_{g}(\theta)$, was assumed to belong to the same probability distribution family as the user-specified prior, $\pi(\theta)$, but with potentially different hyperparameter values. For each outcome type and each scenario defined by $\pi(\theta)$, we simulated 50,000 trials. In the $k$th simulated trial, we first sampled $\theta^{k}$ from $\pi_{g}(\theta)$, then generated interim data ${\cal D}^{k}_{n}=\{X^{k}_{1},\cdots,X^{k}_{n}\}$ from $f(X_{i}|\theta^{k})$, e.g., $\mbox{Bernoulli}(\theta^{k})$ for binary endpoints or $N(\theta^{k},\sigma^{2})$ for normal endpoints. We implemented two interim analyses at the 40th and 70th patients, and a final analysis at the 100th patient, i.e., for n=40, 70n=40,\ 70, and $100$, the trial was stopped and claim that the treatment is effective if $\mbox{Pr}(\theta>\theta_{0}|{\cal D}^{k}_{n})>C$.
@@
As a benchmark, we also evaluated the same performance metrics for trials conducted without interim analyses (i.e., fixed design). In these trials, the decision of claiming that the treatment is effective was made at the end of the trial based on the criterion $\mbox{Pr}(\theta>\theta_{0}|{\cal D}^{k}_{N})>C$. The probability cutoff $C$ was calibrated to control the pFDR at 5% under the scenario where $\pi(\theta)=\pi_{g}(\theta)$. This same cutoff was then applied to designs with and without interim analyses; therefore, any differences observed can be directly attributed to the inclusion of interim analyses. It is important to note that we did not re-calibrate the decision cutoff $C$ for each individual scenario. As a result, in scenarios where $\pi(\theta)\neq\pi_{g}(\theta)$, the pFDR and other metrics for the fixed design may deviate from the nominal level of 0.05. However, this does not pose a concern for our investigation, as the primary objective is to compare results with and without including interim analyses to assess their impact on operating characteristics, rather than to evaluate the absolute operating characteristics of a design across different scenarios.
@@
Below are the simulation parameters for single-arm trials; those for RCTs are provided in Supplementary Materials. For binary endpoint, we set the response rate of the (historical) control as $\theta_{0}=0.6$ and assumed the true response rate of the experimental arm generated from $\pi_{g}(\theta)=$ Beta(3, 3). We evaluated three classes of user-specified prior $\pi(\theta)$: (i) neutral priors centered at $\theta_{0}=0.6$; (ii) pessimistic priors centered at $0.5$; and (iii) optimistic priors centered at $0.7$. Within each class, we examined three priors with increasing spread (i.e., decreasing informativeness). Specifically, for $\pi(\theta)$ centered at 0.6, we considered Beta(18, 12), Beta(3.6, 2.4), and Beta(0.06, 0.04). For $\pi(\theta)$ centered at 0.5, we considered Beta(15, 15), Beta(3, 3), and Beta(0.05, 0.05). Notably, Beta(3, 3) matches $\pi_{g}(\theta)$. And for $\pi(\theta)$ centered at 0.7, we considered Beta(21, 9), Beta(4.2, 1.8), and Beta(0.07, 0.03). The probability cutoff was calibrated as previously described, resulting in $C=0.689$. Due to the discrete nature of the observed data, it was not always possible to find a cutoff yielding an exact 5% pFDR. This value of $C$ was selected to achieve an pFDR closest to, but not exceeding, 5%.
@@
For the normal outcome with $\sigma^{2}=1$, we assumed $\pi_{g}(\theta)=N(0,1)$ and set $\theta_{0}=0.25$. The neutral priors $\pi(\theta)$ centered at 0.25 included $N(0.25,0.2),N(0.25,1)$, and $N(0.25,1000)$. Optimistic and pessimistic priors were centered at 0.5 and 0, respectively. The probability cutoff was calibrated and set to $C=0.55$. For the normal outcome with unknown variance, we used the conjugate prior: $\pi_{g}(\theta|\sigma^{2})=N(0,\sigma^{2}/5)$ and $\pi_{g}(\sigma^{2})=\mbox{Inv-}\chi^{2}(5,40)$, with $\theta_{0}=0.25$. Priors $\pi(\theta,\sigma^{2})$ were constructed with $\theta$ centered at 0.25 (neutral), 0 (pessimistic), and 0.5 (optimistic), as detailed in Table 4. The probability cutoff was set at $C=0.63$.
@@
For the time-to-event outcomes, we assumed a 12-month follow-up period with patient accrual following a Poisson process at a rate of six patients per month. We set $\theta_{0}=8$ and used $\pi_{g}(\theta)=\mbox{Gamma}(12,2)$ and $\pi_{g}(\kappa)=\mbox{Gamma}(8,2)$. We evaluated three $\pi(\theta,\kappa)$ priors with $\theta$ centered at 8 (neutral), 6 (pessimistic), and 10 (optimistic), each with varying precision (see Table 5). The cutoff was calibrated and set to $C=0.64$ to control the pFDR at 5% in the absence of an interim analysis.
@@
The performance metrics were calculated as follows. pFDR was calculated as the proportion of trials in which the treatment is not effective (i.e., $\theta\leq\theta_{0}$) among those claiming that the treatment is effective, across the 50,000 simulated trials. FDR was then obtained by multiplying the pFDR by the proportion of trials claiming that the treatment is effective. Bias was determined as the average difference $\hat{\theta}^{k}-\theta^{k}$, and MSE was computed as the mean of $(\hat{\theta}^{k}-\theta^{k})^{2}$, where $\hat{\theta}^{k}$ denotes the posterior mean of $\theta$ in the $k$th simulated trial. Coverage was defined as the percentage of trials in which $\theta^{k}$ fell within the 95% posterior credible interval for $\theta$ in the $k$th trial. FPR was computed as the percentage of trials claiming that the treatment is effective among those in which the treatment is actually not effective with $\theta\leq\theta_{0}$.
@@ 3.2 Results
Table 2 presents results for single-arm trials with a binary endpoint, using the fixed design without interim analyses as the reference. Unless otherwise noted, the discussion focuses on comparing designs with interim analyses to the corresponding fixed design. This allows us to isolate and assess the impact of conducting multiple interim looks, rather than comparing absolute operating characteristics across different scenarios. Our findings are consistent with the theoretical investigation in Section 2.3, indicating that interim analyses can have a substantial effect on the operating characteristics of the design.
In the ideal scenario where the user-specified prior matches the data-generating prior (i.e., $\pi(\theta)=\pi_{g}(\theta)=$Beta(3, 3)), adding interim analyses does not affect the coverage rates of 95% credible intervals, or the bias. However, incorporating interim analyses lead to inflation in pFDR, FDR, ATIE, and MSE. For example, incorporating interim analyses increased the pFDR from 0.046 to 0.095, FDR from 0.012 to 0.028, ATIE from 0.018 to 0.041, and MSE from 0.002 to 0.003.
@@
Tables 3 and 4 present analogous results for normal endpoints with known and unknown variances, and Table 5 shows findings for time-to-event endpoints. Across all endpoint types, patterns were broadly consistent with those described above for the binary case. Results for RCTs are generally similar to those of single-arm trials; see Tables S1-S4 in Supplementary Materials for more details.
@@ 3.3 Sensitivity analyses
We examined the effect of altering the frequency of interim analyses. For binary and normal outcomes, where patient-level data are typically observed promptly, we employed a fully sequential design in place of the original two interim analyses at the 40th and 70th patients. In this design, early stopping was assessed after each patient beginning with the 20th, i.e., for $n=20,\cdots,N$. For time-to-event endpoints, where outcomes are event-driven rather than patient-driven, interim analyses were conducted every 5 patients starting from the 20th, to allow time for meaningful accumulation of event information. Results for each outcome type under these settings are shown in Tables S5-S7 and were broadly consistent with those observed under the original design.
We further evaluated the performance metrics under alternative values of the probability that a treatment is truly effective, denoted by $\mbox{Pr}_{g}({\rm effective})=\Pr_{\theta\sim\pi_{g}}(\theta>\theta_{0})$, as determined by the data-generating prior distribution of $\theta$. In the context of drug development, $\mbox{Pr}_{g}({\rm effective})$ represents the proportion of all investigational drugs that are genuinely effective. For example, under a normal outcome with known variance 1 and an equivalence margin $\delta=0$, the prior $\pi_{g}(\theta)=N(0,1)$ yields $\mbox{Pr}_{g}({\rm effective})=0.4$. We also considered a more concentrated prior, $\pi_{g}(\theta)=N(0,0.09)$, which reduces $\mbox{Pr}_{g}({\rm effective})$ to 0.2. The corresponding results for continuous, binary, and time-to-event outcomes in single-arm trials, presented in Tables S8 - S10 of the Supplementary Materials, exhibited patterns consistent with those observed under the original prior specification.
For the normal outcome with unknown variance, the scale parameter $\sigma_{0}^{2}$ in $\pi(\sigma^{2})$ was originally set to 40, matching that in $\pi_{g}(\sigma^{2})$. In sensitivity analyses, we varied $\sigma^{2}_{0}$ to 20 and 60. The resulting metrics, shown in Tables S11 and S12, remained consistent with those from the main setting, further supporting the robustness of the findings.
@@ 4 Conclusion
We have conducted a comprehensive investigation into the impact of repeated Bayesian interim analyses on key design metrics, including the bias and MSE of the posterior mean, the coverage probability of posterior credible intervals, FDR/pFDR, and ATIE. Our results show that the claim—that Bayesian interim analyses with superiority stopping are free from multiplicity concerns and may be performed at any time without adjustment—holds for the posterior mean and credible-interval coverage only when the user-specified prior coincides exactly with the true data-generating prior, an assumption that is rarely, if even, satisfied in practice. Moreover, we show that even under this ideal scenario, repeated interim analyses can still systematically distort FDR, pFDR, and ATIE. More importantly, when the user-specified prior differs from the true data-generating prior, which is typically the case in real-world applications, all of these metrics can be adversely affected. These theoretical findings are supported by extensive simulation studies, which confirm that repeated interim analyses can substantially alter the operating characteristics of Bayesian trials, especially when vague or noninformative priors are used. This is particularly concerning because such priors are commonly employed in practice.
@@
Table 1: Impact of interim superiority stopping on design performance metrics, relative to the fixed design without interim analyses, assuming the same decision rule for declaring effectiveness.
Metric Including interims when $\pi=\pi_{g}$ Justification Including interims when $\pi\neq\pi_{g}$ Justification
Posterior mean bias Maintains unbiased Follows from the law of total expectation under the correct prior. Biased Prior misspecification combined with selective stopping distorts the posterior mean.
FDR/pFDR Increased Multiple interim looks increase the chance of stopping on spurious signals when the treatment actually is not effective. Often further increased Prior misspecification distorts posterior probabilities, often worsening FDR and pFDR.
ATIE Increased More looks at the data raise the probability of false positives. Often further increased Same mechanism, compounded by posterior miscalibration.
Not necessarily increased, as it depends on whether $\pi$ is optimistic or pessimistic on efficacy relative to $\pi_{g}$.
Table 2: Single-arm trials with binary endpoints: pFDR, FDR, bias, MSE, coverage of posterior credible intervals, and ATIE, evaluated with or without interim stopping, under various user-specified priors with $\theta_{0}=0.6$ and data-generating prior $\pi_{g}(\theta)=\mbox{Beta}(3,3)$.
$\pi(\theta)$ Interims pFDR FDR Bias ($\times 10^{-3}$) MSE Coverage ATIE
Beta(18, 12) no 0.060 0.017 23.0 0.004 0.793 0.025
yes 0.123 0.039 19.1 0.005 0.777 0.057
Beta(3.6, 2.4) no 0.060 0.017 5.4 0.002 0.949 0.025
yes 0.123 0.039 7.7 0.003 0.948 0.057
Beta(0.06, 0.04) no 0.060 0.017 -0.2 0.002 0.948 0.025
yes 0.131 0.042 5.4 0.003 0.944 0.062
Beta(15, 15) no 0.024 0.006 -0.1 0.003 0.822 0.008
yes 0.038 0.009 -7.40 0.0 0.798 0.014
Beta(3, 3) no 0.046 0.012 -0.3 0.002 0.950 0.018
yes 0.095 0.028 -0.1 0.003 0.949 0.041
Beta(0.05, 0.05) no 0.060 0.017 -0.3 0.002 0.948 0.025
yes 0.131 0.042 5.3 0.003 0.944 0.062
Beta(21, 9) no 0.122 0.040 46.0 0.005 0.713 0.059
yes 0.265 0.109 50.9 0.007 0.692 0.160
Beta(4.2, 1.8) no 0.079 0.023 11.1 0.002 0.943 0.034
yes 0.166 0.057 16.0 0.003 0.940 0.084
Beta(0.07, 0.03) no 0.060 0.017 -0.1 0.002 0.948 0.025
yes 0.131 0.042 5.6 0.003 0.944 0.062
User-specified prior matches the data-generating prior
Table 3: Single-arm trials with normal endpoints with a known variance of 1: pFDR, FDR, bias, MSE, coverage of posterior credible intervals, and ATIE, evaluated with or without interim stopping, under various user-specified priors with $\theta_{0}=0.25$ and data-generating prior $\pi_{g}(\theta)=N(0,1)$.
$\pi(\theta)$ Interims pFDR FDR Bias ($\times 10^{-3}$) MSE Coverage ATIE
N(0.25, 0.2) no 0.051 0.021 12.2 0.012 0.925 0.035
yes 0.098 0.043 -0.8 0.020 0.915 0.072
N(0.25, 1) no 0.051 0.021 2.7 0.010 0.949 0.035
yes 0.098 0.043 4.2 0.016 0.949 0.072
N(0.25, 1000) no 0.051 0.021 0.2 0.010 0.950 0.035
yes 0.098 0.043 6.0 0.017 0.950 0.072
N(0, 0.2) no 0.045 0.018 0.3 0.011 0.927 0.030
yes 0.084 0.036 -19.3 0.021 0.912 0.060
N(0, 1) no 0.050 0.020 0.2 0.010 0.950 0.034
yes 0.095 0.041 0.3 0.016 0.949 0.069
N(0, 1000) no 0.051 0.021 0.2 0.010 0.950 0.035
yes 0.098 0.043 6.0 0.017 0.950 0.072
N(0.5, 0.2) no 0.058 0.024 24.1 0.012 0.919 0.040
yes 0.113 0.051 18.1 0.020 0.911 0.085
N(0.5, 1) no 0.053 0.022 5.1 0.010 0.949 0.036
yes 0.100 0.044 8.2 0.016 0.949 0.074
N(0.5, 1000) no 0.051 0.021 0.2 0.010 0.950 0.035
yes 0.098 0.043 6.0 0.017 0.950 0.072
User-specified prior matches the data-generating prior
Table 4: Single-arm trials with normal endpoints with an unknown variance: pFDR, FDR, bias, MSE, coverage of posterior credible intervals, and ATIE, evaluated with or without interim stopping, under various user-specified priors with $\theta_{0}=0.25$ and data-generating prior $\pi_{g}(\theta|\sigma^{2})=N(0,\sigma^{2}/5)$ and $\pi_{g}(\sigma^{2})=\mbox{Inv-}\chi^{2}(5,40)$.
$\pi(\theta)$ Interims pFDR FDR Bias ($\times 10^{-3}$) MSE Coverage ATIE
$\theta|\sigma^{2}\sim N(0.25,\sigma^{2}/20)$ no 0.050 0.022 39.4 0.837 0.895 0.041
$\sigma^{2}\sim Inv-\chi^{2}(20,40)$ yes 0.094 0.045 -103.4 1.465 0.872 0.085
$\theta|\sigma^{2}\sim N(0.25,\sigma^{2}/5)$ no 0.051 0.023 11.6 0.633 0.950 0.043
$\sigma^{2}\sim Inv-\chi^{2}(5,40)$ yes 0.099 0.048 15.4 1.017 0.950 0.090
$\theta|\sigma^{2}\sim N(0.25,\sigma^{2}/0.1)$ no 0.052 0.023 0.7 0.664 0.950 0.043
$\sigma^{2}\sim Inv-\chi^{2}(0.1,40)$ yes 0.100 0.049 90.5 1.119 0.949 0.092
$\theta|\sigma^{2}\sim N(0,\sigma^{2}/20)$ no 0.046 0.020 -2.3 0.836 0.896 0.038
$\sigma^{2}\sim Inv-\chi^{2}(20,40)$ yes 0.084 0.039 -162.1 1.494 0.871 0.074
$\theta|\sigma^{2}\sim N(0,\sigma^{2}/5)^{*}$ no 0.050 0.022 -0.4 0.633 0.950 0.041
$\sigma^{2}\sim Inv-\chi^{2}(5,40)^{*}$ yes 0.095 0.046 -3.6 1.013 0.950 0.086
$\theta|\sigma^{2}\sim N(0,\sigma^{2}/0.1)$ no 0.052 0.023 0.4 0.664 0.950 0.043
$\sigma^{2}\sim Inv-\chi^{2}(0.1,40)$ yes 0.100 0.049 90.0 1.119 0.949 0.092
$\theta|\sigma^{2}\sim N(0.5,\sigma^{2}/20)$ no 0.055 0.025 81.0 0.842 0.893 0.046
$\sigma^{2}\sim Inv-\chi^{2}(20,40)$ yes 0.106 0.052 -43.2 1.447 0.873 0.098
$\theta|\sigma^{2}\sim N(0.5,\sigma^{2}/5)$ no 0.053 0.023 23.5 0.634 0.950 0.044
$\sigma^{2}\sim Inv-\chi^{2}(5,40)$ yes 0.102 0.049 34.4 1.020 0.950 0.093
$\theta|\sigma^{2}\sim N(0.5,\sigma^{2}/0.1)$ no 0.052 0.023 0.9 0.664 0.950 0.043
$\sigma^{2}\sim Inv-\chi^{2}(0.1,40)$ yes 0.100 0.049 90.8 1.119 0.949 0.092
User-specified prior matches the data-generating prior
Table 5: Single-arm trials with time-to-event endpoints: pFDR, FDR, bias, MSE, coverage of posterior credible intervals, and ATIE, evaluated with or without interim stopping, under various user-specified priors with $\theta_{0}=8$ and data-generating prior $\pi_{g}(\theta)=\mbox{Gamma}(12,2)$ and $\pi_{g}(\kappa)=\mbox{Gamma}(8,2)$.
$\pi(\theta)$ Interims pFDR FDR Bias ($\times 10^{-3}$) MSE Coverage ATIE
$\theta\sim\mbox{Gamma}(128,16)$ no 0.072 0.009 166.5 0.121 0.739 0.010
$\kappa\sim\mbox{Gamma}(32,8)$ yes 0.160 0.022 134.5 0.215 0.725 0.025
$\theta\sim\mbox{Gamma}(21.33,2.67)$ no 0.061 0.007 36.0 0.051 0.942 0.008
$\kappa\sim\mbox{Gamma}(8,2)$ yes 0.211 0.031 59.0 0.155 0.939 0.036
$\theta\sim\mbox{Gamma}(0.64,0.08)$ no 0.058 0.007 -0.6 0.051 0.948 0.008
$\kappa\sim\mbox{Gamma}(0.16,0.04)$ yes 0.357 0.066 538.7 3.985 0.936 0.075
$\theta\sim\mbox{Gamma}(72,12)$ no 0.017 0.002 -15.8 0.072 0.892 0.002
$\kappa\sim\mbox{Gamma}(32,8)$ yes 0.022 0.002 -53.3 0.153 0.877 0.003
$\theta\sim\mbox{Gamma}(12,2)^{*}$ no 0.050 0.006 -1.6 0.049 0.949 0.007
$\kappa\sim\mbox{Gamma}(8,2)^{*}$ yes 0.116 0.015 -5.7 0.104 0.948 0.017
$\theta\sim\mbox{Gamma}(0.36,0.06)$ no 0.059 0.007 -1.8 0.051 0.949 0.008
$\kappa\sim\mbox{Gamma}(0.16,0.04)$ yes 0.351 0.064 572.8 4.867 0.936 0.073
$\theta\sim\mbox{Gamma}(200,20)$ no 0.235 0.038 428.4 0.362 0.420 0.043
$\kappa\sim\mbox{Gamma}(32,8)$ yes 0.859 0.770 2759.4 10.062 0.122 0.882
$\theta\sim\mbox{Gamma}(33.33,3.33)$ no 0.082 0.010 91.9 0.064 0.914 0.012
$\kappa\sim\mbox{Gamma}(8,2)$ yes 0.571 0.164 476.7 1.090 0.858 0.188
$\theta\sim\mbox{Gamma}(1,0.1)$ no 0.059 0.007 1.2 0.051 0.949 0.008
$\kappa\sim\mbox{Gamma}(0.16,0.04)$ yes 0.368 0.069 550.3 3.772 0.937 0.079
User-specified prior matches the data-generating prior